the length of a rectangle is 16 feet more than twice its length. if the perimeter is 104 feet, find the length of the rectangle
Answer by elima(1433)

Lenth = 2w+16
formula;
104=2(2w+16+w)
104=2(3w+16)
104=6w+32
104-32=6w
72=6w
12=w
length=2(12)+16
length=24+16
Length=40
Hope you understand
=)